# Provable Ordering and Continuity in Vision-Language Pretraining for Generalizable Embodied Agents

### v1

@@ Appendix A Proofs
All time steps $t\in[1,T]$ are covered by at least one sampled interval, ensuring the entire video sequence satisfies the Brownian Bridge regularization. Now, let $\mathbf{v}_{k},\mathbf{v}_{l}\in\mathbb{R}^{d}$ be arbitrary embeddings, not necessarily the endpoints $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ of a sampled interval. These embeddings fall within the union $\mathfrak{U}$ of all sampled local intervals. Without loss of generality, here we can identify the interval $[n(i),n(j)]\in\mathfrak{U}$ from the union containing $\mathbf{v}_{k}$ and $\mathbf{v}_{l}$.
@@
Appendix B Pre-training Details
Following (Ma et al., 2023a; Li et al., 2024), we use a modified ResNet-50  (He et al., 2016) from CLIP  (Radford et al., 2021) for the vision encoder and a CLIP transformer for the language encoder. We initialize our model with CLIP and train them on EPIC-KITCHEN-100  (Damen et al., 2018, 2020). The training hyperparameters used during the pre-training are listed in Table 3. For $\mathcal{L}_{BB}$, due to the large number of video frames, we apply a logarithmic scaling to the variance term. The training was conducted on two NVIDIA A800 GPUs, and the process took approximately 30 hours.
Table 3: Hyper-parameters for pre-training.
Config value
Training epochs 1000
Optimizer Adam
Learning rate $1\times 10^{-5}$
Batch size 128
Sampled frames per video 10
Weight decay 0.001
Optimizer momentum 0.9, 0.999
Data augmentation RandomCropResize
Appendix C Evaluation Details
C.1 Simulation Environments
We follow  (Nair et al., 2022) for the specific simulation environment setup and code details.
Franka Kitchen.
The Franka Kitchen environment  (Gupta et al., 2019; Fu et al., 2020) is based on the 9 degrees of freedom Franka robot. The Franka robot is placed in a kitchen environment containing several common household items: a microwave, a kettle, an overhead light, cabinets, and an oven. Following  (Nair et al., 2022), the Franka Kitchen environments used in this paper are modified from their original design. Specifically, we introduce additional randomization to the scene by randomly altering the kitchen’s position between episodes. This modification makes the tasks significantly more challenging in terms of both perception and control.
Metaworld.
The Metaworld environment  (Yu et al., 2019) is an open-source simulated benchmark for robot learning. In our settings, the target object position is randomized between episodes in all tasks.
We present the specific language instructions for each tasks in Table 4.
Environment ID Language Instruction
kitchen_micro_open-v3 open microwave
kitchen_sdoor_open-v3 slide cabinet
kitchen_ldoor_open-v3 open left door
kitchen_knob1_on-v3 turn on stove
kitchen_light_on-v3 switch on light
hammer-v2-goal-observable hammer nail
button-press-topdown-v2-goal-observable press button
assembly-v2-goal-observable assemble the ring onto peg
C.2 Language-Conditioned Behavior Cloning Hyperparameters
We present the LCBC imitation learning hyperparameters in Table 5. For each distinct evaluation task, we perform policy evaluation every 1,000 gradient steps by running 50 rollouts and computing their average success rate. Over a total of 10,000 gradient steps, we conduct this evaluation 10 times. The highest success rate among these 10 evaluations is reported as the final result. To ensure robustness, we average the results across two different camera viewpoints and three independent random seeds.
Table 5: Hyper-parameters for LCBC.
Franka Kitchen Metaworld
MLP achitecture [256,256] [256,256]
Non-linear activation ReLU ReLU
Optimizer Adam Adam
Gradient Steps 10K 10K
Learning rate $1\times 10^{-3}$ $1\times 10^{-3}$
Batch size 32 32
Horizon 50 100
Proprioception 9 4
C.3 Language-Conditioned Behavior Cloning Results
Table 6: LCBC results when dataset size$=5$ on Franka Kitchen.
CLIP 38.7 2.0 3.0 7.0 7.7 11.7
R3M 68.7 18.3 7.7 19.3 29.0 28.6
LIV 55.0 6.0 7.0 13.0 22.0 20.6
DecisionNCE 59.3 9.7 7.0 26.3 24.3 25.3
AcTOL w/o BB 71.5 11.5 10.5 23.5 47.0 32.8
AcTOL 85.5 20.0 18.3 24.7 62.3 42.6
Table 7: LCBC results when dataset size$=15$ on Franka Kitchen.
CLIP 71.0 8.0 15.7 14.7 28.0 27.5
R3M 81.0 31.0 22.0 19.3 57.7 42.2
LIV 85.0 19.0 28.3 29.7 51.7 42.7
DecisionNCE 92.0 18.7 27.0 33.3 45.0 43.2
AcTOL w/o BB 84.5 29.5 29.5 54.0 73.5 54.2
AcTOL 99.5 37.5 37.0 53.5 81.5 61.8
Table 8: LCBC results when dataset size$=25$ on Franka Kitchen.
CLIP 66.3 8.7 18.7 23.7 38.7 31.2
R3M 84.7 35.3 40.0 34.0 61.7 51.1
LIV 91.7 26.0 35.0 45.3 61.7 51.9
DecisionNCE 91.7 27.0 37.0 47.3 51.3 50.9
AcTOL w/o BB 92.0 37.0 40.0 57.0 78.0 60.8
AcTOL 100.0 37.0 42.5 62.5 81.0 64.6
Table 9: LCBC results when dataset size$=5$ on Metaworld.
Method Assembly Pick bin Press button Hammer Average
CLIP 48.3 35.3 34.3 51.2 42.3
R3M 63.5 33.3 27.3 63.2 46.8
LIV 61.8 32.3 32.7 61.0 47.0
DecisionNCE 54.0 31.0 27.7 65.7 44.6
AcTOL w/o BB 66.8 39.0 20.7 74.7 50.3
AcTOL 62.8 41.0 42.0 69.5 53.8
Table 10: LCBC results when dataset size$=15$ on Metaworld.
Method Assembly Pick bin Press button Hammer Average
CLIP 73.0 40.3 52.0 76.0 60.3
R3M 80.7 17.0 45.0 83.3 56.5
LIV 84.3 37.0 54.7 81.3 64.3
DecisionNCE 73.3 36.7 43.3 83.0 59.1
AcTOL w/o BB 94.0 50.3 48.3 90.7 70.8
AcTOL 82.5 64.5 65.5 84.0 74.1
Table 11: LCBC results when dataset size$=25$ on Metaworld.
Method Assembly Pick bin Press button Hammer Average
CLIP 69.3 36.0 66.0 78.8 62.5
R3M 87.7 14.7 48.3 89.7 60.1
LIV 87.3 23.7 66.0 89.7 66.7
DecisionNCE 85.7 47.0 58.0 88.3 69.8
AcTOL w/o BB 93.7 51.7 55.0 93.0 73.3
AcTOL 93.5 66.0 76.5 88.5 81.1
Figure 6: Reward plots for exemplar robot action videos.
C.4 Language-Conditioned Visual Reward Results
As shown in Figure 6, we present more visualizations of Language-Conditioned Visual Reward on real-world robot manipulation videos from  (Bahl et al., 2022). In Figure 6(a), the robot performs two consecutive and opposing actions. Our method effectively identifies the action boundaries and generates the correct reward sequence, increasing first and then decreasing, in alignment with the given instructions. In Figures 6(b)-(d), where the robot performs a single action, the robot initially moves slowly as it searches for the target. Correspondingly, the reward grows gradually. Once the robot interacts with the object and completes the task, our method captures the distinct semantic changes in the action, leading to a rapid reward increase. In Figures 6(e)-(f), we test two complex actions and instructions to explore the limits of our method. In Figure 6(e), the model is required to accurately distinguish between the blue and red cups to complete the task. In Figure 6(f), the model needs to differentiate the orientation and face values of two dice. These scenarios impose high demands on the model’s visual and semantic understanding. Our method successfully produces the correct rewards in both tasks, showcasing its potential for application in real-world, complex scenarios.

### v3

@@ C.2 Proofs of Theorem 2
All time steps $t\in[1,T]$ are covered by at least one sampled interval, ensuring the entire video sequence satisfies the Brownian Bridge regularization. Now, let $\mathbf{v}_{k},\mathbf{v}_{l}\in\mathbb{R}^{d}$ be arbitrary embeddings, not necessarily the endpoints $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$ of a sampled interval. These embeddings fall within the union $\mathfrak{U}$ of all sampled local intervals. Without loss of generality, here we can identify the interval $n(i),n(j)]\in\mathfrak{U}$ from the union containing $\mathbf{v}_{k}$ and $\mathbf{v}_{l}$.
@@ C.3 Proofs of Theorem 3
Appendix D Broader Impacts
We introduce Action Temporal Coherence Learning (AcTOL), a vision-language pretraining framework aimed at improving the generalization capabilities of embodied agents in a variety of manipulation tasks. By learning from large-scale human action videos, AcTOL helps agents acquire temporally consistent representations aligned with natural language, which can support more flexible and data-efficient robotic learning. However, some potential risks should be acknowledged. If AcTOL is trained on video data that contains societal biases or stereotypes, those patterns may be reflected in the model’s behavior. For instance, if certain groups or actions are underrepresented or portrayed inaccurately, the resulting agents could behave in ways that are inappropriate or unreliable in diverse real-world settings. While these challenges are common across many data-driven systems in robotics and vision-language learning, we believe future work should explore strategies such as dataset auditing, fairness-aware training, and improved transparency to support more responsible and robust deployment.